OpenStudy (anonymous):

Determine if (-3/10 *-2) is a solution to 10x-9=15

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please show work!

OpenStudy (anonymous):

in order for that to be right x would have to equal. \[x=\frac{12}{5}\]

OpenStudy (anonymous):

10x-9=15 10x=24 x=24/10=12/5

OpenStudy (anonymous):

Thanks girl!!

OpenStudy (anonymous):

Beehive was right thank them not me

OpenStudy (anonymous):

thanks Beehive as well!!